# ProxNF: Neural Field Proximal Training for High-Resolution 4D Dynamic Image Reconstruction

### v1

@@ V-B Study 2: Dynamic reconstruction
In this study, the proposed ProxNF method was assessed and compared against a well-established spatiotemporal image reconstruction method with nuclear norm regularization (STIR-NN). Specifically, a 2D cross-sectional slice and 3D volume extracted from the dynamic tumor perfusion phantom employed in Study 1 were virtually imaged using the stylized photoacoustic tomography system described in Section IV-B.
TABLE III: Study 2: Dynamic reconstruction. Number of parameters, RRMSE, SSIM, ROI RRMSE, ROI SSIM, and LAC RRMSE achieved by the proposed ProxNF method and the reference STIR-NN method in two (2D) and three (3D) spatial dimensions.
@@
The number of parameters required to represent the reconstructed estimate of the object and the corresponding accuracy metrics—RRMSE and SSIM for the entire image, the RRMSE and SSIM for the region of interest with the inserted lesion, and the lesion activity curve (LAC) RRMSE—are reported in Table III. As displayed by RRMSE and SSIM over the entire image and the region of interest, the ProxNF reconstruction was more accurate compared to the STIR-NN method. Additionally, the ProxNF reconstruction led to a more accurate estimation of the lesion activity, as indicated by the LAC RRMSE and illustrate in Fig. 7. Furthermore, the ProxNF reconstruction required less than a tenth of the parameters used by the STIR-NN. Selected frames of these reconstructions and the original object are shown in Fig. 8.
@@
The number of parameters and accuracy metrics for the ProxNF and STIR-NN methods are reported in Table III. The reconstructed time activities for the inserted lesion are shown in Fig. 9. Selected frames of these reconstructions and the original object are shown in Fig. 10. Note that the ProxNF reconstructions is more accurate in terms of RRMSE and SSIM, both over the entire image and the region of interests, and provides a more accurate time activity curve.
@@
Figure 7: Study 2: Average lesion activity of the dynamic object in two spatial dimensions (blue), nuclear norm reconstruction (green), and neural field reconstruction (red). Note how ProxNF is more accurate in the pre- and post-injection phases.
@@
Figure 8: Study 2: Spatiotemporal image reconstruction in two spatial dimensions. Dynamic object (top) and its estimates produced by STIR-NN (middle), and ProxNF (bottom) are shown at selected imaging frames. The dynamic range of the grayscale image is $[10^{-04},0.06]$. A logarithmic scale was used in the color map to increase visualization contrast at depth. An animation of each reconstructed estimate is available in the supplemental multimedia materials (Video 3). ProxNF can produce an object estimate with comparable accuracy to that produced by STIR-NN, while reducing the memory requirements of one order of magnitude.
@@
Figure 9: Study 2: Average lesion activity of true object in three-spatial dimensions(blue) and neural field reconstruction (red). The neural field representation underestimates the true lesion activity but matches its behavior.
@@
Figure 10: Study 2: Spatiotemporal image reconstruction in three spatial dimensions. Paraview [65] was used for volume rendering. Dynamic object (top) and its estimates produced by STIR-NN (middle), and ProxNF (bottom) are shown at selected imaging frames. Dynamic range of the rendering is [0-0.05] using a linear scale. An animation of each reconstruction from this study is available in the supplemental multimedia materials (Video 4). The proposed method can better capture the dynamic changes in the object compared to the reference method (STIR-NN). Furthermore, as shown in Video 4, ProxNF can resolve the respiratory motion of the numerical phantom while the reconstructed estimate produced reference method only captures a time-averaged dynamic over the respiratory cycle.
VI Conclusion
@@
This works proposes ProxNF, a proximal splitting method for high-resolution, high-dimensional image reconstruction using a neural field (NF) representation. NFs are a special class of neural networks that provide accurate representations of continuous objects and can be utilized for memory-efficient image reconstruction. NFs have demonstrated benefits in dynamic image reconstruction, including automatic exploitation of spatiotemporal redundancies[20, 17, 13, 14]. The proposed ProxNF method increases the scalability of NF-based image reconstruction approaches enabling high-resolution, long time horizon volumetric dynamic imaging. ProxNF achieves this by separating computations involving updates of the NF parameters from those involving the imaging operator, thereby reducing the number of evaluations of the imaging operator and allowing for efficient training of NF parameters. The efficacy of the proposed partition-of-unity-based NF architecture and ProxNF method were illustrated in two numerical studies motivated by dynamic contrast-enhanced photoacoustic computed tomography imaging applications.
@@
The first study showed that the proposed partition of unity network (POUnet) architecture is able to accurately represent a dynamic images in two and three spatial dimensions, achieving an accuracy comparable to that yielded by a space-time semiseparable approximation using a similar number of parameters. The study also showed that the proposed architecture has better representation properties than the original POUnet architecture in [13].
@@
The second study considered dynamic image reconstruction problems in two and three spatial dimensions. A numerical tumor perfusion phantom was virtually imaged with a limited number of views per imaging frame. This study demonstrated that the NF architecture trained using ProxNF outperforms a well-established image reconstruction method utilizing nuclear norm regularization in terms of RRMSE, SSIM, while using a smaller number of parameters to represent the sought-after image. Furthermore, the NF method yielded a more accurate estimate of the lesion activity curve compared to that achieved by the method using nuclear norm regularization.
@@
In summary, this work established the feasibility of NF representations for dynamic image reconstruction of high-resolution, high-dimensional images. ProxNF, the proposed proximal splitting method, helps overcome the computational shortcomings of previous approaches to spatiotemporal image reconstruction using NFs [13]. A key advantage of the proposed method is that it is agnostic of the chosen neural representation. The plug-and-play formulation of ProxNF allows to seamlessly replace the POUnet architecture used in this work with any other state-of-the-art neural field architecture[19, 17, 18, 20, 21, 13]. The primary limitation of this work is the use of numerical phantoms, future studies should investigate the application of the proposed method to in-vivo (clinical or preclinical) data. Additionally, another limitation is the relatively long computational time, however, training the NF can be accelerated by use of modern machine learning software[66]. Future work may extend the proposed method to 5D (three spatial dimensions, one time dimension, and one wavelength dimensions) quantitative PACT, in which dynamic changes in the object optical properties are estimated from multispectral PACT data by use of a non-linear multiphysics imaging operator coupling photon transport and wave propagation.
@@ Appendix A Code and Data Availability
Appendix B Supplementary Multimedia Material
Supplementary multimedia materials include the following. Video 1 depicts the results of the representation study in two spatial dimensions displayed in Fig. 4 alongside the time activity curves for the inserted lesion. Video 2 depicts the representation results in three spatial dimensions displayed in Fig. 6 alongside the time activity curves and the underlying anatomy. Video 3 depicts the results of the dynamic image reconstruction study in two spatial dimensions displayed in Fig. 8 alongside the time activity curves for the inserted lesion. Video 4 depicts the dynamic image reconstruction results in three spatial dimensions displayed in Fig. 10 alongside the time activity curves and the underlying anatomy.

### v3

@@ V-B Study 2: Dynamic reconstruction
In this study, the proposed ProxNF method was assessed and compared against a well-established spatiotemporal image reconstruction method with nuclear norm regularization (STIR-NN) and, for the problem in two spatial dimensions, with the gradient-based NF method (GradNF) in [13]. Specifically, a 2D cross-sectional slice and 3D volume extracted from the dynamic tumor perfusion phantom employed in Study 1 were virtually imaged using the stylized PACT system described in Section IV-B.
TABLE III: Study 2: Dynamic reconstruction. Number of parameters, RRMSE, SSIM, ROI RRMSE, ROI SSIM, and LAC RRMSE achieved by the proposed ProxNF method, gradient-based NF method (GradNF 2D), and the reference STIR-NN method in two (2D) and three (3D) spatial dimensions.
@@
TABLE IV: Study 2: Dynamic reconstruction. Number of evaluations of the (randomly sub-sampled) imaging operator (# evals), batch size (J), and wall-time (W-time).
# evals J W-time
STIR-NN 2D 140 1,296 17 CPU minutes
GradNF 2D 128,000 324 4.0 GPU hours
ProxNF 2D 1,000 324 1.3 GPU hours
STIR-NN 3D 1272 324 110 CPU hours
ProxNF 3D 800 324 200 GPU hours
@@
The number of parameters required to represent the reconstructed estimate of the object and the corresponding accuracy metrics—RRMSE and SSIM for the entire image, the RRMSE and SSIM for the region of interest with the inserted lesion, and LAC RRMSE—are reported in Table III. The ProxNF and GradNF reconstructions achieved similar accuracy and outperformed the STIR-NN reconstruction with respect to all evaluation metrics. Furthermore, the ProxNF and GradNF reconstructions required less than a tenth of the parameters used by the STIR-NN. The LAC and selected frames of the object and its estimates are shown in Fig. 7 and 8, respectively. Table IV summarizes the computational cost of the three methods in terms of number of evaluations of the imaging operator and wall-time. Note that, while STIR-NN accessed all $K=1,296$ imaging frames at each iteration to evaluate the imaging operator, ProxNF and GradNF stochastically subsampled the imaging operator using $J=324$ randomly selected frames at each iteration. While slower than STIR-NN, ProxNF provides significant acceleration ($\sim$128X less evaluations of the imaging operator, $\sim$3X reduction in wall-time) compared to GradNF.
Additionally, manifold discovery and analysis (MDA) [77] was performed on the learned partition network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ in Eq. (26) at each ProxNF iteration. MDA provides a two-dimensional, easy-to-visualize representation of a neural network outputs and inner layers. This analysis shows that $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ quickly stabilizes to a fixed point, thus providing empirical evidence of the ProxNF convergence properties (see Remark 2). A video showing the evolution of the MDA representation of $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ at each ProxNF iteration is available in the supplementary materials (Video 4).
@@
The number of parameters and accuracy metrics for the ProxNF and STIR-NN methods are reported in Table III. The reconstructed LACs are shown in Fig. 9. Selected frames of these reconstructions and the original object are shown in Fig. 10. Note that the ProxNF reconstruction is more accurate than the STIR-NN reconstruction with respect to all metrics and requires less evaluations of the imaging operator (c.f. Table IV).
@@
Figure 7: Study 2a: Dynamic reconstruction in two spatial dimensions. LAC of the object (blue), and its estimates obtained by STIR-NN (green), GradNF (purple), and ProxNF (red). Note how the NF estimates are more accurate than the STIR-NN one in the pre- and post-injection phases.
@@
Figure 8: Study 2a: Dynamic reconstruction in two spatial dimensions. Object (top row) and its estimates produced by STIR-NN (second row), GradNF (third row), and ProxNF (last row) are shown at selected imaging frames. The dynamic range of the grayscale image is $[10^{-4},0.06]$. A logarithmic scale was used to increase visualization contrast at depth. An animation of each reconstructed estimate is available in the supplemental multimedia materials (Video 3). ProxNF can produce an object estimate with comparable accuracy to that produced by STIR-NN while reducing the memory requirements of one order of magnitude.
@@
Figure 9: Study 2b: Dynamic reconstruction in three spatial dimensions. LAC of the object (blue) and its estimates produced by STIR-NN (green) and ProxNF (red). The NF representation underestimates the true LAC less severely than STIR-NN.
@@
Figure 10: Study 2b: Dynamic reconstruction in three spatial dimensions. ParaView [76] was used for volume rendering. Dynamic object (top) and its estimates produced by STIR-NN (middle), and ProxNF (bottom) are shown at selected imaging frames. Dynamic range of the rendering is $[0,0.05]$ using a linear scale. An animation of each reconstruction from this study is available in the supplemental multimedia materials (Video 5). The proposed method can better capture the dynamic changes in the object compared to the reference method (STIR-NN). Furthermore, as shown in Video 4, ProxNF can resolve the respiratory motion of the numerical phantom while the reconstructed estimate produced reference method only captures a time-averaged dynamic over the respiratory cycle.
VI Application to in-vivo DCE PACT imaging of tumor perfusion
The proposed ProxNF method was also demonstrated using retrospective in-vivo DCE PACT imaging data of tumor perfusion in small animal models. Animal studies were conducted in compliance of the current regulations and standards of the National Institutes of Health and approved by the Institutional Animal Care and Use Committee of MD Anderson Cancer Center. In this study, a single cross-sectional slice of a mouse torso bearing a pancreatic cancer xenograft flank tumor is dynamically imaged at a single light wavelength.
VI-A Experimental setup and data acquisition
Measurement data were acquired utilizing the commercially available inVision PACT animal scanner (iThera Medical; Munich Germany). This system employs a toroidally focused, ring-shaped transducer array with 256 detection elements covering 270 around the sample. Transducers’ center frequency is 5 MHz and sampling rate 50 MHz. Together with a near-infrared laser (680–980 nm) with a 10 Hz repetition rate, this setup enables cross-sectional tomographic imaging of a small animal torso with an in-plane resolution of 150$\sim$180 $\mu$m. In the experiment, data were continuously acquired at 8 different wavelengths, 6 averages per wavelength before, during and after the injection of the contrast agent. The total duration of the scan was $T=16$ minutes, resulting in 200 imaging frames with a temporal resolution of 4.8 seconds.
VI-B Image reconstruction studies
An image reconstruction study was performed utilizing undersampled measurements at a single wavelength (800 nm). Specifically, at each imaging frame, only data collected by 64 transducer elements (8 equispaced groups of 8 contiguous elements each) was employed. The subset of active transducer elements was cyclically rotated for each frame. Image reconstruction using the proposed ProxNF method and the STIR-NN method was performed on a $256\times 256$ spatial grid (square field of view with size 36 mm) and 200 imaging frames.
The NF implemented the same network architecture described in the numerical studies. The nuclear norm reconstruction utilized a maximum rank of 20.
A reference image was estimated utilizing measurements from all the 256 transducer elements at all time frames by use of the fast spatiotemporal image reconstruction method in [69]. Based on Morozov discrepancy principle, the rank of the truncated singular value decomposition of the data matrix $\boldsymbol{d}$ was set to 20 in the estimation of the reference image. Accuracy of the ProxNF and STIR-NN reconstructions were then assessed using RRMSE and SSIM, evaluated both on the entire image and on a ROI containing the lesion, with respect to the reference image. Image quality was also assessed by analyzing the estimated LAC, which is the time-varying spatially-averaged signal over a volume of interest (the segmented lesion).
VI-C Results and analysis
The number of parameters required to represent the reconstructed image and the corresponding accuracy metrics—RRMSE and SSIM for the entire image, the RRMSE and SSIM for the ROI containing the lesion, and the LAC RRMSE—are reported in Table V. The ProxNF reconstruction was more accurate compared to the STIR-NN method with respect to all metrics. Furthermore, the ProxNF reconstruction required half the number of the parameters used by the STIR-NN. The LAC and selected frames of the reference image and the estimates produced by the ProxNF and STIR-NN methods are shown in Figs. 11 and 12, respectively.
TABLE V: In-Vivo Experiment: Number of parameters, RRMSE, SSIM, ROI RRMSE, ROI SSIM, and LAC RRMSE achieved by the ProxNF and STIR-NN methods with respect to the reference image
# of RRMSE SSIM ROI ROI LAC
Params RRMSE SSIM RRMSE
STIR-NN 1.31 M $0.2441$ $0.9617$ $0.2706$ $0.8379$ $0.0908$
ProxNF 730 K $0.1876$ $0.9767$ $0.2449$ $0.8745$ $0.0846$
Figure 11: In-Vivo Experiment: LAC of the reference image (blue), nuclear norm reconstruction (green), and NF reconstruction trained using ProxNF (red).
Figure 12: In-Vivo Experiment: Reference image (top) reconstructed utilizing full measurements, STIR-NN (middle), and ProxNF (bottom) reconstructed images from undersampled measurements are shown at selected imaging frames. The dynamic range of the grayscale image is $[10^{-4},0.6]$. A logarithmic scale is employed to improve image contrast at depth. An animation of each reconstructed estimate is available in the supplemental multimedia materials (Video 6). ProxNF produces an object estimate with improved accuracy compared to the STIR-NN reconstruction, while halving the memory requirements.
VII Conclusion
@@
This work proposes ProxNF, a proximal splitting method for high-resolution, high-dimensional image reconstruction using a neural field (NF) representation. NFs are a special class of neural networks that provide accurate representations of continuous objects and can be utilized for memory-efficient image reconstruction. NFs have demonstrated benefits for several medical imaging applications [37], including dynamic imaging[20, 17, 13, 14]. The proposed ProxNF method increases the scalability of NF-based image reconstruction approaches enabling high-resolution, long-time horizon volumetric dynamic imaging. ProxNF achieves this by separating computations involving updates of the NF parameters from those involving the imaging operator, thereby reducing the number of evaluations of the imaging operator and allowing for efficient training of NF parameters. The efficacy of the proposed partition-of-unity-based NF architecture and ProxNF method were illustrated in two numerical studies and in-vivo dynamic contrast-enhanced photoacoustic computed tomography (DCE PACT) imaging of tumor perfusion in small animal models.
@@
The first numerical study showed that the proposed partition of unity network (POUnet) architecture is able to accurately represent a dynamic image in two and three spatial dimensions, achieving an accuracy comparable to that yielded by a space-time semiseparable approximation using a similar number of parameters.
@@
The second numerical study considered dynamic image reconstruction problems in two and three spatial dimensions. A numerical tumor perfusion phantom was virtually imaged with a limited number of views per imaging frame. This study demonstrated that the NF architecture trained using ProxNF outperforms a well-established image reconstruction method utilizing nuclear norm regularization (STIR-NN) in terms of RRMSE and SSIM while using a smaller number of parameters to represent the sought-after image. Furthermore, the NF method yielded a more accurate estimate of the lesion activity curve compared to that achieved by the method using nuclear norm regularization.
In the in-vivo DCE PACT study, an axial cross-section of a mouse was estimated utilizing undersampled measurements at each imaging frame. The accuracy of the proposed method was assessed utilizing a reference image reconstructed from the fully sampled measurements, and compared to that achieved by STIR-NN. In this application, ProxNF achieved a higher accuracy, as well as a 50% reduction in the number of parameters, compared to STIR-NN.
@@
In summary, this work established the feasibility of NF representations for dynamic image reconstruction of high-resolution, high-dimensional images. ProxNF, the proposed proximal splitting method, helps overcome the computational shortcomings of previous approaches to spatiotemporal image reconstruction using NFs [13]. A key advantage of the proposed method is that it is agnostic of the chosen neural representation. The plug-and-play formulation of ProxNF allows to seamlessly replace the POUnet architecture used in this work with any other state-of-the-art NF architecture[19, 17, 18, 20, 21, 13]. Additionally, ProxNF could easily be generalized to other approaches for representing images with neural networks, such as deep image priors [26], and still provide a separation of concerns between optimizing network parameters and evaluations of the imaging operator.
The primary limitation of this work is the relatively long computational time compared to well-established image reconstruction method utilizing nuclear norm regularization. However, the use of high performance deep-learning backends, such as JAX[78] and improved NF embedding algorithms can potentially mitigate this issue. Future work may extend the proposed method to 5D (three spatial dimensions, one time dimension, and one wavelength dimensions) quantitative PACT, in which dynamic changes in an object’s chromophore concentrations are estimated from multispectral PACT data by use of a non-linear multiphysics imaging operator coupling photon transport and wave propagation.